40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 13026     Entry time: Thu Jun 1 00:10:15 2017     In reply to: 13019
Author: gautam 
Type: Update 
Category: General 
Subject: Coil driver boards reinstalled 

[Koji, Gautam]

We tried to debug the mysterious sudden failure of ASS - here is a summary of what we did tonight. These are just notes for now, so I don't forget tomorrow.

What are the problems/symptoms?

  • After re-installing the coil driver electronics, the ASS loops do not appear to converge - one or more loops seem to run away to the point we lose the lock.
  • For the Y-arm dithers, the previously nominal ITM PIT and YAW oscillator amplitudes (of ~1000cts each) now appears far too large (the fuzz on the Y arm transmission increases by x3 as viewed on StripTool).
  • The convergence problem exists for the X arm alignment servos too.

What are the (known) changes since the servos were last working?

  • Gain of x3 on the de-whitening boards for ITMX, ITMY, BS and SRM have been replaced with gain x1. But I had measurements for all transfer functions (De-White board input to De-White Board outputs) before and after this change, so I compensated by adding a filter of gain ~x3 to all the coil filter banks for these optics (the exact value was the ratio of the DC gain of the transfer functions before/after).
  • The ETMY Oplev has been replaced. I walked over to the endtable and there doesn't seem to be any obvious clipping of either the Oplev beam or the IR transmission.

Hypotheses plus checks (indented bullets) to test them:

  1. The actuation on the ITMs are ~x10 times stronger now (for reasons unknown).
    • I locked the Y-arm and drove a line in the channels C1:SUS-ETMY_LSC_EXC and C1:SUS-ITMY_LSC_EXC at ~100Hz and ~30Hz, (one optic at one frequency at a time), and looked at the response in the LSC control signal. The peaks at both frequencies for the ITMs and ETMs were within a factor of ~2. Seems reasonable.
    • We further checked by driving lines in C1:SUS-ETMY_ASCPIT_EXC and C1:SUS-ITMY_ASCPIT_EXC (and also the corresponding YAW channels), and looked at peak heights at the drive frequencies in the OL control signal spectra - the peak heights matched up well in both the ITM and ETM spectra (the drive was in the same number of counts).

      So it doesn't look like there is any strange actuation imbalance between the ITM and ETM as a result of the recent electronics work, which makes sense as the other control loops acting on the suspensions (local damping, Oplevs etc seem to work fine). 
  2. The way the dither servo is set up for the Y-arm, the tip-tilts are used to set the input axis to the cavity axis, while actuation to the ITM and ETM takes care of the spot centering. The problem lies with one of these subsystems.
    • We tried disabling the ASS servo inputs to all the spot-centering loops - but even with just actuation on the TTs, the arm transmission isn't maximized.
    • We tried the other combination - disable actuation path to TTs, leave those to ITM and ETM on - same result, but the divergence is much faster (lock lost within a couple of seconds, large offsets appear in the ETM_PIT_L / ETM_YAW_L error signals.
    • Tried turning on loops one at a time - but still the arm transmission isn't maximized.
  3. Something is funny with the IR transmon QPD / ETMY Oplev.
    • I quickly measured Oplev PIT and YAW OLTFs, they seem normal with upper UGFs around 5Hz and phase margins of ~30 degrees.
    • We had no success using either of the two available Transmon QPDs
    • Looking at the QPD quadrants, the alignment isn't stellar but we get roughly the same number of counts on all quadrants, and the spot isn't drastically misaligned in either PIT or YAW.

For whatever reasons, it appears that dithering the cavity mirrors at frequencies with amplitudes that worked ~3 weeks ago is no longer giving us the correct error signals for dither alignment. We are out of ideas for tonight, TBC tomorrow...

 

ELOG V3.1.3-